OpenStudy (anonymous):

Help please! Gallium has two naturally occurring isotopes (Ga-69 and Ga-71) and has an atomic mass of 69.723 amu. The mass of Ga-71 is 70.925 amu, and its natural abundance is 39.892%. Calculate the atomic mass (in amu) and natural abundance (in %) of the second isotope.

OpenStudy (anonymous):

(atomic weight of one isotope)(x) + (atomic weight of other isotope)(1-x) = the atomic mass Use algebra to find x. Don't forget to use the distributive property. Example: Cl-35 weighs 34.969 amu Cl-37 weighs 36.966 amu (34.969)(x) + (36.966)(1-x) = 35.453 (on the table for Chlorine) 34.969 x + 36.966 - 36.966 x = 35.453 -1.997 x = -1.513 x = 0.75764 so Cl-35 abundance is 75.764% and Cl-37 is 24.236 %

OpenStudy (anonymous):

you got it??

OpenStudy (anonymous):

hold on..i am trying to process what you wrote ha

OpenStudy (anonymous):

hmm this is actually kinda of confusing for me..I'll show you how I did it. 60.108% Ga-69 : 0.60108 x ______amu= 39.892 % Ga-71: 0.39892x 70.925 amu= 28.293 amu 100-39.892 = 60.1085% Ga-69 69.723-28.293 = 41.4796 am Ga-69 Is this correct?

OpenStudy (anonymous):

tada..ur right

OpenStudy (anonymous):

still wait lemme check

OpenStudy (anonymous):

Some of my classmates divide the 41.4796 by 0.60108? are they correct for doing that

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

hmm yeah me too...okay thanks!

OpenStudy (anonymous):

yor welcomed